# RIDE: Difficulty Evolving Perturbation with Item Response Theory for Mathematical Reasoning

### v1

@@ 6 Conclusion
In this paper, we propose a question rewriting framework based on difficulty evolution. The method estimates question difficulty using IRT, then trains a difficulty ranker to provide a difficulty reward signal, guiding reinforcement learning for the rewriting model. Experiments show that our method causes significant performance drops in $26$ LLMs, validating its effectiveness and robustness in adversarial perturbation. Our method outperforms traditional rule-based perturbations and can be applied for training data augmentation. Further analyses confirm its effectiveness, offering a new perspective on robust evaluation from the view of “question difficulty”.
@@
All our source code will be submitted in the Supplementary Materials, along with an explanation to facilitate reviewers in reproducing our work. The hyperparameter setup is introduced in Appendix B.2. We will also open source our mathematical question rewriting models RIDE-4B and RIDE-8B, our perturbation benchmarks RIDE-AMC and RIDE-AIME, and our augmented training data RIDE-DeepMath.
Appendix A The Use of LLMs
In the experimental phase, LLMs are mainly used in the following aspects:
  • As base models: For example, our question rewriting model RIDE-8B is training upon Qwen3-8B.
  • As data filter: We employ GPT-5 and DeepSeek-V3.1 to filter out questions that are unsolvable or have incorrect answers.
  • As evaluators: During the reinforcement learning stage, we use GPT-5-mini to provide correctness rewards. Besides, GPT-5 serves as a teacher model to judge the win rate between two rewritten questions.
  • As evaluated models: We evaluate the feasibility of our proposed method for difficulty-evolution robust perturbations on $26$ LLMs.
In the non-experimental phase, LLMs are applied, including:
  • Code assistance: Using GPT-5 for environment setup, code debugging, and partial script development. The LLMs-related code has been rigorously reviewed by humans.
  • File processing: Using GPT-5 to handle files in formats such as JSON, and summarize some results in training logs.
  • Paper writing support: Leveraging GPT-5 for grammar checking and polishing to improve accuracy and quality of academic writing.
Appendix B Experimental Setup
B.1 Implementation Details
We use the API interface provided by OpenAI to call models and obtain the response matrix of LLMs (for the GLM series, we use the ZHIPU interface). The size of the response matrix is 3535×$2000$. We perform IRT parameter estimation through py-irt (Lalor_2023). The statistics of each student model’s performance is shown in Figure 6, along with the accuracy and estimated ability $\theta$ are shown in Table 5.
@@ B.1 Implementation Details
In the SFT stage, we use the LLaMA-Factory (zheng-etal-2024-llamafactory) framework for training, with a dataset of $2,000$ rewritten mathematical questions distilled from GPT-5-mini. In the reinforcement learning stage, we use the TRL (vonwerra2022trl) framework for training, with $3,000$ questions extracted from the DeepMath dataset (he2025deepmath103klargescalechallengingdecontaminated). In our experiments, model performance improved after benchmark perturbations in a small number of cases (Table B.1), which is more common with Qwen3 in thinking mode, which further confirms its strong robustness and strong reasoning ability.
Model Params AIME-24 RIDE-AIME PDR (↓%) AMC-23 RIDE-AMC PDR (↓%)
[…]
Grok-3 - 60.00 63.33 $\uparrow$5.55 92.50 87.50 5.41
[…]
Qwen3-14B 14B 73.33 76.67 $\uparrow$4.55 80.00 70.00 12.50
Qwen3-235B-A22B 235B 90.00 93.33 $\uparrow$3.70 97.50 95.00 2.56
Table 6: Cases of performance improvement on perturbed benchmarks.
B.2 Hyperparameter Setup
The hyperparameters used in our work are summarized in Table B.2. For the IRT parameter estimation and the training of the difficulty ranker, we perform a greedy search to select the optimal group of hyperparameters based on evaluation metrics. For the SFT and reinforcement learning stages, the hyperparameters are chosen by balancing model performance with the available computational resources.
@@ B.2 Hyperparameter Setup
Deepspeed (deepspeed) Deepspeed Stage 3
[…]
Device Device used for SFT 4 * NVIDIA A800 80GB PCIe
@@
B.3 Prompt Strategy
@@
For constructing the response matrix and evaluating model performance on both the original benchmark and the perturbed benchmark, we use the same zero-shot prompt to instruct models to answer mathematical questions (Figure 7), requiring only that answers follow a specific format for verification. In practice, we find that even without explicitly telling the model to “think step by step”, all evaluated models still produce chain-of-thought outputs.
@@
Figure 8 shows the prompt we use to guide the model in rewriting mathematical questions during the SFT and reinforcement learning stages. We do not impose a single, detailed rule. Instead, we briefly outline several ways to increase difficulty and allow the model to improvise, making the trajectory of difficulty evolution unpredictable.
@@
Figure 9 presents the prompt we use to instruct GPT-5 to choose the one that better meets the requirements from two rewritten questions, enabling us to compute the win rate of our method against the rule-based approach. We specify in detail that the rewritten question must address covering problem completeness, clarity, conceptual consistency, structural perturbations, and the avoidance of reasoning shortcuts.
Figure 7: Prompt for mathematical question answering.
Figure 8: Prompt for rewriting mathematical questions.
Figure 9: Prompt for judging which rewritten question is better.
Appendix C Effect of IRT models and difficulty ranker.
@@
As described in Section 5.4, we conducted additional experiments to verify the effectiveness of the choice of the IRT model and data augmentation methods. The results are shown in Table 8. We consider two priors provided by py-irt: a vague prior that exerts minimal influence so the likelihood largely drives estimation, and a hierarchical prior that treats parameters as draws from a shared group-level distribution with hyperpriors, enabling partial pooling to stabilize estimates and shrink extremes, especially with sparse data. For evaluation, in addition to AUC-ROC for discrimination, we also report the Brier Score, the mean squared difference between predicted probabilities and observed outcomes (lower is better), which captures overall probabilistic accuracy and calibration and thus complements AUC-ROC’s ranking focus.
@@
LR Regression BGE-M3 (chen2024bgem3embeddingmultilingualmultifunctionality) 1.1380
@@
Longformer (beltagy2020longformerlongdocumenttransformer) Regression Qwen3-8B 1.0746

### v2

@@ 7 Conclusion
In this paper, we propose a question rewriting framework based on difficulty evolution. The method estimates question difficulty using IRT, then trains a difficulty ranker to provide a difficulty reward, guiding reinforcement learning for the rewriting model. Experiments show that our method causes performance drops in $26$ LLMs, validating its effectiveness and robustness in adversarial perturbation. Our method outperforms traditional rule-based perturbations and can be applied for training data augmentation. Further analyses confirm its effectiveness, offering a new perspective on robust evaluation from the view of “question difficulty”.
Impact Statements
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@
All our source code will be submitted in the Supplementary Materials, along with an explanation to facilitate reviewers in reproducing our work. The hyperparameter setup is introduced in Appendix A.2. We will also open source our mathematical question rewriting models RIDE-4B and RIDE-8B, our perturbation benchmarks RIDE-AMC and RIDE-AIME, and our augmented training data RIDE-DeepMath.
Appendix A Experimental Setup
A.1 Implementation Details
We use the API interface provided by OpenAI to call models and obtain the response matrix of LLMs (for the GLM series, we use the ZHIPU interface). The size of the response matrix is 3535×$2000$. We perform IRT parameter estimation through py-irt (Lalor and Rodriguez, 2023). The statistics of each student model’s performance is shown in Figure 5, along with the accuracy and estimated ability $\theta$ are shown in Table 5. The results indicate that within the same model series (e.g., Qwen2.5 or Qwen3), model parameter size is positively correlated with capability. However, this relationship is not numerically linear. Additionally, frontier reasoning models generally exhibit higher capability scores.
@@ A.1 Implementation Details
In the SFT stage, we use the LLaMA-Factory (Zheng et al., 2024) framework for training, with a dataset of $2,000$ rewritten mathematical questions distilled from GPT-5-mini. In the reinforcement learning stage, we use the TRL (von Werra et al., 2020) framework for training, with $3,000$ questions extracted from the DeepMath dataset (He et al., 2025).
A.2 Hyperparameter Setup
The hyperparameters used in our work are summarized in Table 6. For the IRT parameter estimation and the training of the difficulty ranker, we perform a greedy search to select the optimal group of hyperparameters based on evaluation metrics. For the SFT and reinforcement learning stages, the hyperparameters are chosen by balancing model performance with the available computational resources.
@@ A.2 Hyperparameter Setup
Deepspeed (Rasley et al., 2020) Deepspeed Stage 3
Reinforcement Learning
Device Device used for RL 4 * NVIDIA A800 80GB PCIe
@@
A.3 Prompt Strategy
(a) Prompt for mathematical question answering.
(b) Prompt for rewriting mathematical questions.
(c) Prompt for judging which rewritten question is better.
Figure 6: Overview of prompts.
@@ A.3 Prompt Strategy
For constructing the response matrix and evaluating model performance on both the original benchmark and the perturbed benchmark, we use the same zero-shot prompt to instruct models to answer mathematical questions (Figure 6(a)), requiring only that answers follow a specific format for verification. In practice, we find that even without explicitly telling the model to “think step by step”, all evaluated models still produce chain-of-thought outputs.
@@
Figure 6(b) shows the prompt we use to guide the model in rewriting mathematical questions during the SFT and reinforcement learning stages. We do not impose a single, detailed rule. Instead, we briefly outline several ways to increase difficulty and allow the model to improvise, making the trajectory of difficulty evolution unpredictable.
@@
Figure 6(c) presents the prompt we use to instruct GPT-5 to choose the one that better meets the requirements from two rewritten questions, enabling us to compute the win rate of our method against the rule-based approach. We specify in detail that the rewritten question must address covering question completeness, clarity, conceptual consistency, structural perturbations, and the avoidance of reasoning shortcuts.
A.4 Rewrite Strategy
We distill six core rewrite strategies from the rewritten questions: (1) Wording and Paraphrasing Modification, (2) Distracting Info, (3) Numerical Substitution, (4) Add Extra Steps, (5) Change Constraints, and (6) Change Target. The detailed descriptions are shown in Table 7. To determine the strategies involved in each question, we employed a model-based annotation process. We fed each question to both GPT-5 and DeepSeek-V3.2 three times each, for a total of six annotations. A specific strategy was only recorded as being present for a given question if it was selected by a majority (i.e., at least three out of six) of the model runs.
Strategy Definition
Wording Rephrases the question using different vocabulary or structures without altering math logic.
Distracting Adds irrelevant information, definitions, or values to the question description.
Numerical Replaces key numerical values in the question with new values.
Extra Steps Introduces new events or steps that must be incorporated into the calculation.
Constraints Adds stricter limitations or alters core rules to change the solution’s domain.
Target Changes the final expression to a more complex or generalized version.
Table 7: Definitions of the Rewrite Strategies.
Appendix B Effect Analysis
B.1 Orthogonality between question difficulty and robustness verification
The verification of model robustness and the evolution of question difficulty can coexist in our method. We define the robustness of LLMs in mathematical reasoning as follows: for a question $x$ and the model’s response correctness $y\in\{0,1\}$, given a semantic-preserving transformation family $\mathcal{T}$, if $f(x)=y$, then for the vast majority of $t\in\mathcal{T}$, $f(t(x))=y$ holds. Figure 8 shows that the semantic similarity between our rewritten and original questions consistently exceeds 80%, confirming that performance degradation does not stem from semantic disparity. Building on this, we decouple question difficulty from robustness verification. We identify ”surface-level modifications” that preserve reasoning depth. An example is shown in Figure 7.
Figure 7: A case study of the rewritten questions with surface-only strategies.
In this example, the evolution of the question is manifested in the formalized mathematical conditions and changes in phrasing, which do not impact the depth of reasoning required, making this a case for verifying model robustness. For example, Qwen2.5-72B answered the original question correctly but failed on the rewritten question. We collected similar surface-only rewrites and categorized them into three difficulty levels—easy, medium, and hard—based on the accuracy rates of $10$ LLMs on the original questions. The questions rewritten using the surface-only strategy maintain a reasoning depth consistent with the original questions, leaving the surface form as the sole variable. If the model’s performance degradation were solely dependent on difficulty, the magnitude of the performance drop on the rewritten questions should vary significantly across the three difficulty levels. Conversely, if question difficulty and robustness verification can be decoupled, the variations in performance decline across the three difficulty levels should be insignificant. We evaluated the performance degradation between the original and rewritten questions across different difficulty levels (Table 8).
Level Original Accuracy Rewritten Accuracy $\Delta$
Easy 80.00 68.57 11.43
Medium 68.57 55.71 12.86
Hard 62.50 43.75 18.75
Table 8: Performance drop comparison across different difficulty levels.
The results indicate that surface-level rewriting induces performance degradation even on easy questions, thereby highlighting the necessity of robustness verification. Moreover, the performance drop across difficulty levels is not significant, with a particularly negligible variation in the magnitude of change between easy and medium questions. We additionally recorded the counts of ”Correct on Original, Incorrect on Rewritten” and ”Incorrect on Original, Correct on Rewritten” for each paired sample in the surface-level rewriting dataset and calculated the p-value using McNemar’s test. The resulting p-value of $2.3847\times 10^{-11}$ is substantially lower than 0.05, revealing a significant asymmetry. This demonstrates that the performance decline originates from the model’s non-robustness rather than changes in difficulty. In summary, through case study and statistical analysis, we decoupled question difficulty from robustness verification and found that the two are orthogonal, allowing them to coexist within our method.
B.2 Effect of IRT models and difficulty ranker.
  • Variational Autoencoder (VAE) method. We mitigate the scarcity of student responses by training a VAE (Kingma and Welling, 2022) on the observed response matrix $\mathbf{Y}$ to learn a low-dimensional manifold of student behavior. During generation, we draw a latent code $\mathbf{z}$ from the Gaussian prior and decode it into per-question correctness probabilities $\pi(\mathbf{z})$, from which Bernoulli draws yield a synthetic student-question response vector with conditional independence across items given $\mathbf{z}$:
$\tilde{\mathbf{Y}}$ $=\begin{bmatrix}\mathbf{Y}\\ \mathbf{R}_{\mathrm{VAE}}\\ \mathbf{R}_{\mathrm{Sampling}}\end{bmatrix}$, (1)
$\mathbf{R}_{\mathrm{VAE}}$ $=[\mathbf{r}_{\mathrm{VAE}}^{1},\dots,\mathbf{r}_{\mathrm{VAE}}^{S}]^{\intercal},$
$\mathbf{R}_{\mathrm{Sampling}}$ $=[\mathbf{r}_{\mathrm{Sampling}}^{1},\dots,\mathbf{r}_{\mathrm{Sampling}}^{T}]^{\intercal}$
$S$ and $T$ denote the number of synthetic students by VAE and sampling, respectively. Then the modeling ability of IRT calibration for question difficulties can be improved. We perform IRT parameter estimation based on the response matrix $\tilde{\mathbf{Y}}$.
@@ B.1 Orthogonality between question difficulty and robustness verification
As described in Section 6, we conducted additional experiments to verify the effectiveness of the choice of the IRT model and data augmentation methods. The results are shown in Table 9. We consider two priors provided by py-irt: a vague prior that exerts minimal influence so the likelihood largely drives estimation, and a hierarchical prior that treats parameters as draws from a shared group-level distribution with hyperpriors, enabling partial pooling to stabilize estimates and shrink extremes, especially with sparse data. For evaluation, in addition to AUC-ROC for discrimination, we also report the Brier Score, the mean squared difference between predicted probabilities and observed outcomes (lower is better), which captures overall probabilistic accuracy and calibration and thus complements AUC-ROC’s ranking focus.
@@
LR Regression BGE-M3 (Chen et al., 2024) 1.1380
@@
Longformer (Beltagy et al., 2020) Regression Qwen3-8B 1.0746
@@
B.3 Statistics of RIDE rewritten model.
Starting from Qwen3-1.7B and Qwen3-4B, we trained RIDE-1.7B and RIDE-4B, and preserved SFT checkpoints (RIDE-1.7B-SFT, RIDE-4B-SFT, RIDE-8B-SFT) to separate SFT from RL effects. The questions rewritten by each model are evaluated in terms of average generation time, average token count, correctness, and semantic similarity. Generation time and token count use the minimum of three runs, correctness is judged by GPT-5 and similarity is measured via text-embedding-3-large. The results are shown in Figure 8. Compared to Qwen3-8B, our SFT and reinforcement learning models produce longer outputs and slower inference, showing that rewrites expand questions, solutions, and answers. Semantic similarity drops slightly, indicating greater diversity while preserving affinity with the original questions. Reinforcement learning improves correctness over both the original and SFT-only models. RIDE-8B and RIDE-4B demonstrate much higher correctness than RIDE-1.7B, indicating that parameter size influences the correctness of generated questions. However, this difference is not significant between the 4B and 8B models, suggesting a limited effect of scale within this range.
Figure 8: Statistics of average token count, average generation time, correctness and semantic similarity across different models.
B.4 Effect of LLM verifier
We use the proprietary GPT-5 series models to provide correctness rewards and as filtering models. To verify that our method can also use low-cost, open-source models, we employ GPT-5-mini and Deepseek-V3.2 (non-thinking mode) as verifiers. We provide $2,000$ math questions (without answers), along with the solutions and answers generated by two test models (Qwen2.5-Math-7B and Gemma3-12B), as input to these verifiers to judge the correctness of the test models’ responses. The results, shown in Table 11, indicate that the verification accuracy of the closed-source model GPT-5-mini reach over 85%. In contrast, the accuracy of DeepSeek-V3.2 was relatively lower, but still over 65%. However, both verifiers achieve a precision of around 90%. This indicates that both models can precisely assign low reward signals to incorrect solutions and are effective for filtering erroneous data. The relatively low recall, however, implies that the verifiers tend to be conservative, flagging an otherwise correct solution and answer as incorrect. This may be because the model’s solution contains minor errors even when the final answer is correct. This situation is acceptable, given that rule-based or model-based rewards in Reinforcement Learning with Verified Rewards (RLVR) also exhibit a high number of false negatives (Xu et al., 2025a). While this situation affects training efficiency, it does not directly lead to training failure, as our verifier assigns low reward signals to incorrect rewrites. Therefore, our method can also use open-source models as verifiers and is not merely a distillation from proprietary models.
Response Source Verifier Precision Recall F1 Accuracy
Qwen2.5-Math-7B GPT-5-mini 97.96 79.17 87.57 85.70
Deepseek-V3.2 89.48 62.19 73.38 71.30
Gemma3-12B GPT-5-mini 98.38 79.71 88.06 86.00
Deepseek-V3.2 93.84 54.09 68.62 67.95
Table 11: Verification performance of GPT-5-mini and Deepseek-V3.2 verifiers on solutions generated by different response models.
Appendix C Item Response Theory Analysis
C.1 Human Difficulty & LLM Difficulty
In human cognition, the difficulty of a math question is often determined by its required reasoning depth. Consequently, researchers have established the AoPS (Art of Problem Solving) standard to measure the difficulty of math competition questions from a human perspective, and methods have emerged that use LLMs in conjunction with the AoPS standard to automatically assign difficulty ratings to math questions  (Gao et al., 2025). The DeepMath dataset, which we utilize in the IRT modeling phase, also applies this standard to provide difficulty levels for its questions. However, we discover a discrepancy between this human-centric standard and actual LLM performance. After grouping the responses of our $35$ models on $2,000$ questions by their given AoPS difficulty levels, we calculate the mean empirical accuracy for each level. The Spearman and Kendall correlation coefficients between the AoPS difficulty and the mean empirical accuracy are found to be -0.5152 and -0.5111, respectively, indicating that while the expected negative correlation exists (i.e., higher human-perceived difficulty corresponds to lower model accuracy), the relationship is not sufficiently strong to consider them equivalent. Therefore, we conclude that a deviation exists between the human cognitive difficulty defined by the AoPS standard and the empirical error rate demonstrated by LLMs. Furthermore, treating lower empirical accuracy directly as higher ”LLM difficulty” introduces a fundamental flaw: empirical accuracy assigns equal weight to the responses of all tested LLMs. We therefore introduce IRT to define difficulty, which allows us to derive a statistically robust difficulty value by incorporating the ability of all participating LLMs into the calculation.
Additionally, we compare the correlation of both the empirical error rate (ungrouped) and IRT difficulty against the human-perceived difficulty, again using Spearman and Kendall coefficients, as shown in Table 12. The results show that both empirical error rate and our IRT difficulty (fitted on 35 model responses) have a similar correlation with the given human difficulty. Our work is primarily focused on this ”LLM difficulty” rather than human-perceived difficulty.
vs Human-Rated Difficulty Spearman Kendall
Empirical Error Rate 0.3362 0.2470
IRT Difficulty 0.3418 0.2468
IRT Difficulty w/o augmentation 0.3364 0.2429
IRT Difficulty (Qwen-only) 0.2839 0.2031
Table 12: Correlation analysis of empirical error rate, IRT difficulty and human-rated difficulty. Spearman & Kendall measure the strength and direction of the monotonic relationship (or rank-order correlation) between two variables.
We also conduct ablation studies by fitting IRT models on the original response matrix without augmentation, and on a matrix using only the Qwen-series models. The results show that the difficulty values derived solely from the Qwen series ($11$ test-takers) correlate poorly with both empirical accuracy and human difficulty, as the limited number of subjects introduces significant bias into the IRT fit. Besides, the original response matrix without augmentation is also less correlated due to data sparsity. This finding validates the effectiveness of our VAE and Sampling data augmentation methods. By learning from the original response matrix and empirical accuracy to generate a large number of distributionally-consistent virtual responses to expand the matrix, our approach enhances the robustness of the IRT fitting process. This prevents parameter estimation bias caused by an insufficient number of subjects.
C.2 Case for IRT Difficulty
As previously discussed, although both empirical error rate and IRT difficulty can be regarded as values for LLM Difficulty, empirical error rate’s drawback is its assumption of equal weighting for all LLMs, leading to a lack of item discrimination. In Figure 9, we present a case study of two questions. Although both questions have the identical empirical accuracy (37.14%), their IRT difficulties differ significantly: Question 1 has an IRT difficulty of 0.599, whereas Question 2 has an IRT difficulty of 0.778. This discrepancy arises because the profile of models that answer each question correctly is different. Specifically, four high-ability models including Deepseek-R1-Distill-Qwen-1.5B, Gemini-2.5-Flash, GLM-Z1-AIRX, and Qwen3-235B-A22B (see ability values in Table 5) answer Question 1 correctly but fail to answer Question 2. Therefore, even with the same empirical accuracy, Question 1 is more likely to be solved by high-ability LLMs, thus receiving a lower IRT difficulty score. This case validates the advantage of using IRT difficulty over simple empirical error rate, as it successfully distinguishes item difficulty based on the ability of the solvers, not just the raw pass count.
Figure 9: A case study of two questions with identical empirical accuracy and different IRT difficulties.
C.3 Distribution Shift
To validate the stability of our IRT parameter estimations, we collect additional responses for the same $2,000$ items using five extra models: Qwen2-72B, Mistral-large, GPT-oss-20B, Qwen3-1.7B, and GPT-4.1-nano. Among these, Qwen2-72B and Qwen3-1.7B are classified as ”lower-performing LLMs” as their accuracies fell below the $67.98\%$ average accuracy of our initial $35$ LLMs. We then conduct two experiments: first, we added only these weaker LLMs to the original $35$-LLM response matrix; second, we added all $5$ new models to the original 35-LLM matrix. We then fit IRT parameters for both augmented matrices and calculate the Spearman and Kendall correlation coefficients against the original IRT difficulty parameters derived from the initial $35$ LLMs. The results (shown in Table 13) indicate that in both scenarios—whether adding only the weaker models or all five additional models—the changes in the Spearman and Kendall correlation coefficients are minimal. This demonstrates the stability of our method, confirming that our IRT parameter estimations are robust and not significantly perturbed by the inclusion of a certain amount of new response data.
vs $35$-LLMs IRT Difficulty Spearman Kendall
$35$-LLMs + lower-performing $2$-LLMs 1.00 0.9968
$35$-LLMs + $5$-LLMs 1.00 0.9954
Table 13: Correlation analysis between the original $35$-LLM fitted IRT difficulty and those fitted with 22 additional lower-performing LLMs or all 55 additional LLMs.